A rectangular plate submerged in water is 5 by 4 m, the 5-m side being horizontal and the 4-m side being vertical. Determine the magnitude of the force on one side of the plate and the depth to its center of pressure if the top edge is (a) at the water surface; (b) 1 m below the water surface; (c) 100 m below the water surface


1
Expert's answer
2021-04-14T07:43:20-0400

For the position of the top surface...

(a) At the water surface, the force on the side at the center of pressure will be


"F=\rho g\frac h2A,\\\space\\
F=1000·9.8·\frac42·5·4=392000\text{ N}."

(b) 1 m below the water surface:


"F=\rho g\frac{(d+h)}{2}A,\\\space\\
F=0.5·1000·9.8·\frac{4+1}{2}·5·4=490000\text{ N}."

(c) 100 m below the water surface:


"F=\rho g\frac{(d+h)}{2}A,\\\space\\
F=0.5·1000·9.8·\frac{4+100}{2}·5·4=10192000\text{ N}."